OpenStudy (butterflyhope):

Simplify: (3x^2 + 5x - 14) + (5x^2 + 4x + 11) Thank you!

OpenStudy (tkhunny):

Have you considered adding the "like" terms?

OpenStudy (popitree):

add the like terms square with square like that then factorize

OpenStudy (anonymous):

why cant ijust give them the answer

OpenStudy (butterflyhope):

Yes, but i'm not very good at it ;-; I thought the answer was -6x^3 + 20x^3.

OpenStudy (anonymous):

lol thats wrong

OpenStudy (anonymous):

\[(3x^2 + 5x - 14) + (5x^2 + 4x + 11)\] \[=3x^2+5x^2+5x+4x-14+11\] has the like terms next to each other

OpenStudy (butterflyhope):

I know @jetfreak3

OpenStudy (anonymous):

then since \(3+5=8\) you have \[3x^2+5x^2=8x^2\]

OpenStudy (anonymous):

just like 3 dollars and 5 dollars is 8 dollars, so \(3x^2+5x^2=8x^2\)

OpenStudy (anonymous):

also \(5x+4x=9x\) for the same reason

OpenStudy (anonymous):

finally \(-14+11=-3\)

OpenStudy (tkhunny):

@jetfreak3 You cannot just give the answer because that would be wrong.

OpenStudy (butterflyhope):

Thank you c: @satellite73

OpenStudy (anonymous):

yw hope it is more or less clear put it all together for your final answer

OpenStudy (butterflyhope):

I got 8x^2 + 9x + -3.

OpenStudy (butterflyhope):

And or - -3

OpenStudy (tkhunny):

Well, never write that. 8x^2 + 9x - 3